# Multi-Object Hallucination in Vision-Language Models

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) have motivated increasing efforts in adapting them for understanding visual semantics, giving rise to a surge of large vision language models (LVLMs) Alayrac et al. (2022); OpenAI (2023); Reid et al. (2024). These models, whether explicitly trained with grounding data (Zhang et al., 2024b) or without (Liu et al., 2023b), demonstrate an impressive grounded understanding of visual entities. Despite their promising performances on various downstream applications (Nasiriany et al., 2024), LVLMs often suffer from object hallucination Rohrbach et al. (2018); Dai et al. (2023b); Li et al. (2023a), where they produce objects not present in a given image.
@@
Although object hallucination was initially observed in image captioning describing multiple objects (Rohrbach et al., 2018), current benchmarks for object hallucination primarily concentrate on the presence of a single object class rather than individual entities. These benchmarks either verify if an object class mentioned in the caption can ground to an object in the image (Rohrbach et al., 2018; Jing et al., 2023), or probe the model about the existence of an object class, sometimes with additional attributes or relations to other objects (Li et al., 2023a; Lovenia et al., 2023). There are, however, two key limitations with these setups as shown by a case study in Figure 1. First, grounding is not simply one-to-one between objects and classes, but a many-to-many mapping between objects and phrases (Kamath et al., 2021; Ma et al., 2023). For instance, “apple” could potentially correspond to multiple referents in Figure 1(c), and the model doesn’t necessarily need to recognize all of them to provide such a response. Therefore, being able to produce an object that exists in an image does not necessarily indicate that the model is free of hallucinations. Second, explicitly instructing the model to recognize multiple objects poses greater challenges compared to simple yes/no inquiries that contain explicit text descriptions for individual objects. For instance, while the model can correctly identify that a whisk is positioned to the left of a knife when “a whisk” is deliberately prompted, as shown in Figure 1(b-d), it may hallucinate a “fork” when directly prompted to recognize both the whisk and the knife (i.e., Figure 1a). This could be due to the common association between knives and forks, which leads to potential hallucinations when models are tasked to focus on multiple objects at the same time. To enable LVLMs to recognize and reason about multiple objects that often occur in realistic visual scenes and to better quantify the complex phenomena we observed, this paper investigates multi-object hallucination, examining how models may misperceive (e.g., by inventing nonexistent objects or becoming distracted) when tasked to focus on multiple objects concurrently, and which factors cause the hallucinations.
@@ 2 Related Work
There is a growing trend to harness and adapt the powerful large language models (LLMs) for multimodal understanding beyond text (Tsimpoukelli et al., 2021; Alayrac et al., 2022; Yu et al., 2024a; Qian et al., 2024). Especially, visual instruction tuning has gained prominence for its competitive performance with a comparatively moderate amount of data and computational resources, leading to a variety of Large Vision-Language Models (LVLMs) (Liu et al., 2023b, 2024; Dai et al., 2023a; Zhu et al., 2023; Gong et al., 2023; Wang et al., 2023d; Ye et al., 2023; Laurençon et al., 2023a). Grounding datasets have been shown to benefit vision-language pre-training (Lu et al., 2022; Li et al., 2022; Ma et al., 2023). Researchers have developed a family of grounded LVLMs focusing on object grounding to bounding box (Pi et al., 2023; Chen et al., 2023; Zhao et al., 2023; Bai et al., 2023; You et al., 2023; Zhang et al., 2024a; Peng et al., 2024) and segmentation masks (Lai et al., 2023; Zhang et al., 2023; Xia et al., 2024; Rasheed et al., 2024; Zhang et al., 2024b). Of the large space of LVLMs, our work is most related to visual prompting (Yang et al., 2023a, b) and object hallucination (Rohrbach et al., 2018; Dai et al., 2023b). The paragraphs below describe the two lines of work in detail.
@@
LVLMs demonstrate their grounded understanding of user-provided visual cues, giving rise to a practical and user-friendly prompting paradigm known as visual prompting (Yang et al., 2023a, b). Early work on visual prompting in vision-language models can date back to tuning-based methods (Bahng et al., 2022; Yao et al., 2024). Recent studies show that LVLMs demonstrate zero-shot understanding of user-provided visual cues (e.g., a red circle) (Shtedritski et al., 2023; Yang et al., 2023b). This observation allows prompting LVLMs by editing images directly in the pixel space, e.g., by adding visual marks or visual text (Li et al., 2023b). Starting from Set-of-Marks (SoM) prompting (Yang et al., 2023a), several training-free methods have been introduced (Lei et al., 2024; Yang et al., 2024; Wan et al., 2024). Recent work further enhances visual prompting by additional visual instruction tuning with diverse visual prompts overlaid on the images (Cai et al., 2024), or explicit visual pointer tokens in the models (Lai et al., 2023; You et al., 2023; Zhang et al., 2024b). We leverage visual prompting to avoid potential ambiguity in textual descriptions, especially when evaluating multiple object hallucinations for objects of the same class.
@@
Benchmark Design Considerations #Test
Multi. Distr. Source Ref. Eval.
CCEval Zhai et al. (2023) Seen Text N 0.1k
GAVIE Liu et al. (2023a) Mixed Text N 1k
FAITHScore Jing et al. (2023) Unseen Text N 2k
HaELM Wang et al. (2023b) Unseen Text N 5k
M-HalDetect Gunjal et al. (2024) Unseen Text H 0.8k
MMHal-Bench Sun et al. (2023) Unseen Text N,H 0.1k
CHAIR Rohrbach et al. (2018) Unseen A 46k
AMBER Wang et al. (2023a) Unseen Text A 1k
CIEM Hu et al. (2023) Unseen Text A 5k
NOPE Lovenia et al. (2023) Unseen A 3k
POPE Li et al. (2023a) Train Unseen A 0.5k
ROPE (Ours) Test Seen & Vis. A 5k
Unseen
@@
Table 1: An overview of object hallucination benchmark setups. We summarize the number of tested images, and if multiple classes and object class distribution (at training and test time) are considered. The image sources include those seen or unseen during instruction tuning. To refer to an object, textual descriptions and visual cues can be adopted. For evaluation, neural models, humans, and automatic pipelines are used.
@@
Despite their promising performance on benchmarks, these models frequently generate objects that do not exist in the provided images, a problem known as object hallucination (Rohrbach et al., 2018; Dai et al., 2023b). Several methods have been suggested to mitigate the object hallucination issue, such as integrating an external object detector (Zhai et al., 2023), applying visually grounded visual instruction tuning (You et al., 2023; Zhang et al., 2024b) or reinforcement learning (Sun et al., 2023; Gunjal et al., 2024), performing iterative refinement (Zhou et al., 2024), and adapting the decoding strategies (Huang et al., 2024). To quantify progress on mitigating them, various benchmarks have been developed and have revealed the prevalence of object hallucination, even in images that are seen during instruction tuning (Zhai et al., 2023; Liu et al., 2023a). We contrast our ROPE benchmark against existing benchmarks and setups in Table 1. ROPE, which is designed for evaluating multi-object hallucination, is distinguished in several ways. First, we deliberately consider the distribution of object classes within a single image at test time. Object hallucination is observed originally in image captioning applications, where multiple objects are described (Rohrbach et al., 2018). While existing research has demonstrated that the object class distribution in the instruction tuning dataset can influence hallucination patterns (Li et al., 2023a; Zhou et al., 2024; Wang et al., 2023a), the impact of object class distribution within an image at test time remains under-explored. Second, current benchmarks concentrate on the presence of an object class or distinguish instances using textual descriptions like attributes, which can still result in ambiguity and multiple referents. We instead leverage the visual referring prompting setups and use visual cues (i.e., marked bounding boxes) to uniquely refer to objects. Finally, our evaluation is automated, without black-box neural models or human evaluators.
3 Recognition-based Object Probing Evaluation (ROPE)
We introduce the Recognition Object Probing Evaluation (ROPE), an automated protocol for assessing LVLMs in multi-object recognition.
@@ 3.1 Task Setup
To avoid ambiguity from multiple candidate referents when using text prompts, ROPE leverages visual prompts to uniquely refer to objects. ROPE tasks LVLMs with selecting the best matching class for multiple objects, as referred to by the visual prompt, from a predefined set of object classes. Specifically, each sample in the ROPE protocol consists of a quadruple $\{\mathcal{I},\mathcal{L},\langle p_{1},\cdots,p_{n}\rangle,\langle o_{1},\cdots,o_{n}\rangle\}$: (1) an image $\mathcal{I}$ consisting of at least $n$ objects; (2) a natural language instruction $\mathcal{L}$ that specifies the recognition task, including $N$ candidate object classes $c_{1},\cdots,c_{N}$; (3) $n$ visual prompts $p_{1},\cdots,p_{n}$, each queries an object in the image; and (4) $n$ object classes $o_{1},\cdots,o_{n}$ as the answers. In this work, we construct a dataset with $N=50$ and $n=5$, i.e., models are tasked with recognizing 5 objects out of 50 candidate object classes.
Figure 3: A heterogeneous ROPE sample tested with Default multi-object query, where each of the 5 objects belongs to different object classes. We label the output class as either correct or hallucinated.
@@
For a fair comparison that accommodates both open-weight and API-based LVLMs, ROPE explicitly instructs models to generate a formatted output of object classes, e.g., obj1:<class1>, ..., obj5:<class5>. This format enables automated evaluation through simple parsing, avoiding black-box neural models or human evaluators With different analytical purposes, we designed 3 types of task prompts for Multi-Object queries, as illustrated in Figure 2 and described as follows.
@@
  • Teacher-Forcing: This setting eliminates cumulative error, allowing the model to condition on the correct previous context when generating object classes, leading to upper bound performance in multi-object recognition. We similarly force the model to follow the provided template and decode only the object tokens for each of the five objects, but we replace the previously generated object tokens with the ground truth. This essentially follows the few-shot in-context learning setting. Teacher forcing benefits models especially when they take shortcuts by repeating the object class list as ordered in the prompt, e.g., LLaVA-7B (Liu et al., 2023b) and Gemini 1.0 Pro (Team et al., 2023) in Figure 3.
For comparison, we also designed task prompts for Single-Object query. We probe the model to recognize the object referred to by the visual prompts one at a time, repeating this as 5 independent and individual prompts. Unlike Default multi-object query, the model only needs to focus on one object, which can be seen as an extension of the POPE (Li et al., 2023a) setup from yes/no polling to classification. We refer to Appendix A.1 for the prompt templates for each type of task prompt.
@@
GPT4V** 45.46 63.12 34.17 N/A N/A 47.34 64.94 35.45
GPT4O** 63.27 80.29 54.47 N/A N/A 63.45 79.84 53.74
    Mechanistically grounded LVLMs take visual prompts by dedicated pointer tokens. We slightly adapt the text prompt and probe the object classes with the highest probabilities. We also apply such probabilistic probing to LLaVA-7B for comparison, as all of three models adopt Vicuna-7B v1.5 (Chiang et al., 2023) as the base LLM. See Appendix A.1 for details.
@@
3.2 Dataset Construction
Data Sources and Curation.
Since our goal is to evaluate and analyze multi-object hallucination, the image data must contain multiple objects of diverse classes with instance-level semantic annotations. We build our dataset upon existing panoptic segmentation datasets, including MSCOCO-Panoptic (Lin et al., 2014; Caesar et al., 2018) and ADE20K (Zhou et al., 2017), to ensure access to all object instances and their semantic classes. We note that one can build a dataset using the ROPE protocol with any dataset containing multiple objects and their bounding boxes, such as Visual Genome (Krishna et al., 2017). We describe the data curation pipeline in Appendix A.1.
Splits by Query Distributions.
As shown in Figure 3and 4, our initial observations indicate that LVLMs are less likely to hallucinate objects when they recognize the same object class multiple times. However, they tend to make more mistakes when all tasked object classes are different or when a new object class is introduced after multiple repeated tasks. We thus vary the distribution of object classes within each image at test time, dividing ROPE into 4 subsets: Homogeneous, Heterogeneous, and Adversarial, In-the-Wild.
Attending to Data Contamination.
While data contamination has been explicitly handled in most of the existing benchmarks, object hallucination has been observed even in images that appear in the instruction tuning dataset, such as Visual Genome (Zhai et al., 2023; Liu et al., 2023a). To evaluate whether multi-object hallucination can be observed in both seen and unseen images, and to critically determine if training on these images helps reduce hallucinations, we explicitly split our dataset into Seen and Unseen based on the original split of the datasets. [^1] We believe this approach is the best practice, but we also acknowledge that the distinction between seen and unseen images may not be strict. Uncurated web images often overlap with public test images, and researchers have no transparent access to the datasets used to train some of these LVLMs unfortunately (Dodge et al., 2021). Depending on the object query distributions (4 splits) and whether the tested image was seen during instruction tuning (2 splits), we divide the test into 8 folders.
4 Experiments and Results
(a) LLaVA-7B.
(b) LLaVA-13B.
(c) LLaVA-34B.

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) have motivated increasing efforts in adapting them for understanding visual semantics, giving rise to a surge of large vision language models (LVLMs) alayrac2022flamingo; gpt4v; reid2024gemini. These models, whether explicitly trained with grounding data (zhang2024groundhog) or without (liu2023visual), demonstrate an impressive grounded understanding of visual entities. This motivates a new prompting paradigm based on user-provided visual cues, referred to as visual prompting (shtedritski2023does; yang2023dawn; li2023vrptest; yang2023set; yang2023set). Despite their promising performances on various downstream applications (nasiriany2024pivot), LVLMs often suffer from object hallucination rohrbach2018object; dai2023plausible; li2023evaluating, where they produce objects not present in a given image.
@@
Although object hallucination was initially observed in image captioning describing multiple objects (rohrbach2018object), current benchmarks for object hallucination primarily concentrate on the presence of a single object class rather than individual entities. These benchmarks either verify if an object class mentioned in the caption can ground to an object in the image (rohrbach2018object; jing2023faithscore), or probe the model about the existence of an object class, sometimes with additional attributes or relations to other objects (li2023evaluating; lovenia2023negative). There are, however, two key limitations with these setups as shown by a case study in Figure 1. First, grounding is not simply one-to-one between objects and classes, but a many-to-many mapping between objects and phrases (kamath2021mdetr; ma2023world). For instance, “apples” could potentially correspond to multiple referents in Figure 1, and the model doesn’t necessarily need to recognize all of them to provide such a response. Therefore, being able to produce an object that exists in an image does not necessarily indicate that the model is free of hallucinations. Second, explicitly instructing the model to recognize multiple objects poses greater challenges compared to simple yes/no inquiries that contain explicit text descriptions for individual objects. For instance, while the model can correctly identify that a whisk is positioned to the left of a knife when “a whisk” is deliberately prompted, as shown in Figure 1(b-d), it may hallucinate a “fork” when directly prompted to recognize both the whisk and the knife (i.e., Figure 1a). This could be due to the common association between knives and forks, which leads to potential hallucinations when models are tasked to focus on multiple objects at the same time. In real-world applications, multi-object querying is crucial for embodied AI tasks. For example, in a cooking scenario, an agent must recognize multiple ingredients and tools simultaneously to be effective. We also present a case study on autonomous driving (see Figure LABEL:fig:single and Figure LABEL:fig:multi in the appendix), demonstrating how common associations between cars, pedestrians and traffic lights could lead to potential hallucinations. In addition, evaluating multiple objects simultaneously, rather than querying each object individually, can significantly save both time and resources. To enable LVLMs to recognize and reason about multiple objects that often occur in realistic visual scenes and to better quantify the complex phenomena we observed, this paper investigates multi-object hallucination, examining how models may misperceive (e.g., by inventing nonexistent objects or becoming distracted) when tasked to focus on multiple objects concurrently, and which factors cause the hallucinations.
@@ 2 Related Work
There is a growing trend to harness and adapt the powerful large language models (LLMs) for multimodal understanding beyond text (tsimpoukelli2021multimodal; alayrac2022flamingo; yu2024crema; qian2024affordancellm). Especially, visual instruction tuning has gained prominence for its competitive performance with a comparatively moderate amount of data and computational resources, leading to a variety of Large Vision-Language Models (LVLMs) (liu2023visual; liu2023improved; dai2023instructblip; zhu2023minigpt; gong2023multimodal; wang2023visionllm; ye2023mplug; laurenccon2023obelics). Grounding datasets have been shown to benefit vision-language pre-training (lu2022unified; li2022grounded; ma2023world). Researchers have developed a family of grounded LVLMs focusing on object grounding to bounding box (pi2023detgpt; chen2023shikra; zhao2023bubogpt; bai2023qwen; you2023ferret; zhang2024ferret; peng2024grounding) and segmentation masks (lai2023lisa; zhang2023llava; xia2023gsva; rasheed2023glamm; zhang2024groundhog). Of the large space of LVLMs, our work is most related to visual prompting (yang2023set; yang2023dawn) and object hallucination (rohrbach2018object; dai2023plausible). The paragraphs below describe the two lines of work in detail.
@@
LVLMs demonstrate their grounded understanding of user-provided visual cues, giving rise to a practical and user-friendly prompting paradigm known as visual prompting (yang2023set; yang2023dawn). Early work on visual prompting in vision-language models can date back to tuning-based methods (bahng2022exploring; yao2024cpt). Recent studies show that LVLMs demonstrate zero-shot understanding of user-provided visual cues (e.g., a red circle) (shtedritski2023does; yang2023dawn). This observation allows prompting LVLMs by editing images directly in the pixel space, e.g., by adding visual marks or visual text (li2023vrptest). Starting from Set-of-Marks (SoM) prompting (yang2023set), several training-free methods have been introduced (lei2024scaffolding; yang2024fine; wan2024contrastive). Recent work further enhances visual prompt understanding by additional visual instruction tuning with diverse visual prompts overlaid on the images (cai2023making), or explicitly represent visual pointer tokens in the models (lai2023lisa; you2023ferret; zhang2024groundhog). We leverage visual prompting to avoid potential ambiguity in textual descriptions, especially when evaluating multiple object hallucinations for objects of the same class.
@@
{wraptable}
@@
!tr0.5 Benchmark Design Considerations #Test Multi. Distr. Source Ref. Eval. CCEval zhai2023halle Seen Text N 0.1k GAVIE liu2023mitigating Mixed Text N 1k FAITHScore jing2023faithscore Unseen Text N 2k HaELM wang2023evaluation Unseen Text N 5k M-HalDetect gunjal2024detecting Unseen Text H 0.8k MMHal-Bench sun2023aligning Unseen Text N,H 0.1k CHAIR rohrbach2018object Unseen A 46k AMBER wang2023llm Unseen Text A 1k CIEM hu2023ciem Unseen Text A 5k NOPE lovenia2023negative Unseen A 3k POPE li2023evaluating Train Unseen A 0.5k ROPE (Ours) Test Seen & Vis. A 5k Unseen An overview of object hallucination benchmarks. For design considerations, we summarize the number of tested images, and if multiple classes and object class distribution (at training and test time) are considered. The image sources include those seen or unseen during instruction tuning. To refer to an object, textual descriptions and visual cues can be adopted. For evaluation, neural models, humans and automatic pipelines are used.
@@
Despite their promising performance on benchmarks, these models frequently generate objects that do not exist in the provided images, a problem known as object hallucination (rohrbach2018object; dai2023plausible). Several methods have been suggested to mitigate the object hallucination issue, such as integrating an external object detector (zhai2023halle), applying visually grounded visual instruction tuning (you2023ferret; zhang2024groundhog) or reinforcement learning (sun2023aligning; gunjal2024detecting), performing iterative refinement (zhou2023analyzing), and adapting the decoding strategies (huang2023opera). To quantify progress on mitigating them, various benchmarks have been developed and have revealed the prevalence of object hallucination, even in images that are seen during instruction tuning (zhai2023halle; liu2023mitigating). We contrast our ROPE benchmark against existing benchmarks and setups in Table 4. ROPE, which is designed for evaluating multi-object hallucination, is distinguished in several ways. First, we deliberately consider the distribution of object classes within a single image at test time. Object hallucination is observed originally in image captioning applications, where multiple objects are described (rohrbach2018object). While existing research has demonstrated that the object class distribution in the instruction tuning dataset can influence hallucination patterns (li2023evaluating; zhou2023analyzing; wang2023llm), the impact of object class distribution within an image at test time remains under-explored. Second, current benchmarks concentrate on the presence of an object class or distinguish instances using textual descriptions like attributes, which can still result in ambiguity and multiple referents. We instead leverage the visual referring prompting setups and use visual cues (i.e., marked bounding boxes) to refer to objects. Finally, our evaluation is automated, without black-box neural models or human evaluators.
3 Recognition-based Object Probing Evaluation
We introduce the Recognition Object Probing Evaluation (ROPE), an automated protocol for assessing LVLMs in multi-object recognition. ROPE specifically measures object hallucination in VLMs within a multi-object setting, examining how models may misperceive (e.g., by inventing nonexistent objects or becoming distracted) when tasked to focus on multiple objects concurrently, and which factors cause the hallucinations.
@@ 3.1 Task Setup
To avoid ambiguity from multiple candidate referents when using text prompts, ROPE leverages visual prompts to uniquely refer to objects. ROPE tasks LVLMs with selecting the best matching class for multiple objects, as referred to by the visual prompt, from a predefined set of object classes. Specifically, each sample in the ROPE protocol consists of a quadruple $\{\mathcal{I},\mathcal{L},\langle p_{1},\cdots,p_{n}\rangle,\langle o_{1},\cdots,o_{n}\rangle\}$: (1) an image $\mathcal{I}$ consisting of at least $n$ objects; (2) a natural language instruction $\mathcal{L}$ that specifies the recognition task, including $N$ candidate object classes $c_{1},\cdots,c_{N}$; (3) $n$ visual prompts $p_{1},\cdots,p_{n}$, each queries an object in the image; and (4) $n$ object classes $o_{1},\cdots,o_{n}$ as the answers. In this work, we construct a dataset with $N=50$ and $n=5$, i.e., models are tasked with recognizing 5 objects out of 50 candidate object classes. Although we use this dataset as an example, ROPE can be applied to any dataset containing multiple objects and their bounding boxes.
@@
For a fair comparison that accommodates both open-weight and API-based LVLMs, ROPE explicitly instructs models to generate a formatted output of object classes, e.g., obj1:<class1>, ..., obj5:<class5> (Figure 2). This format enables automated evaluation through simple parsing. This format enables automated evaluation through simple parsing, avoiding black-box neural models or human evaluators With different analytical purposes, we designed 3 types of task prompts for Multi-Object queries, as illustrated in Figure 7 and described as follows.
@@
  • Teacher-Forcing: This setting eliminates cumulative error, allowing the model to condition on the correct previous context when generating object classes, leading to upper bound performance in multi-object recognition. We similarly force the model to follow the provided template and decode only the object tokens for each of the five objects, but we replace the previously generated object tokens with the ground truth. This essentially follows the few-shot in-context learning setting. Teacher forcing helps especially when LVLMs take shortcuts by repeating the object class list as ordered in the prompt (e.g., LLaVA-7B (liu2023visual) and Gemini 1.0 Pro (team2023gemini) in Figure 2).
For comparison, we also designed task prompts for Single-Object query. We probe the model to recognize the object referred to by the visual prompts one at a time, repeating this as 5 independent and individual prompts. Unlike Default multi-object query, the model only needs to focus on one object, which can be seen as an extension of the POPE (li2023evaluating) setup from yes/no polling to classification. We refer to Appendix A.1 for the prompt templates for each type of task prompt.
Figure 2: A heterogeneous ROPE sample tested with Deafult multi-object query, where each of the 5 objects belongs to different object classes. We label the output class as either correct or hallucinated.
3.2 Dataset Construction
Data Sources and Curation.
Since our goal is to evaluate and analyze multi-object hallucination, the image data must contain multiple objects of diverse classes with instance-level semantic annotations. We build our dataset upon existing panoptic segmentation datasets, including MSCOCO-Panoptic (lin2014microsoft; caesar2018coco) and ADE20K (zhou2017scene), to ensure access to all object instances and their semantic classes. We note that one can build a dataset using the ROPE protocol with any dataset containing multiple objects and their bounding boxes. We describe the data curation pipeline in Appendix A.1.
Splits by Query Distributions.
As shown in Figure 2and 3, our initial observations indicate that LVLMs are less likely to hallucinate objects when they are tasked with recognizing the same object class multiple times. However, they tend to make more mistakes when all tasked object classes are different or when a new object class is introduced after multiple repeated tasks. We thus consider the distribution of object classes within each image at test time, dividing ROPE into 4 subsets: Homogeneous, Heterogeneous, and Adversarial, In-the-Wild.
  • Adversarial: The first 4 tested objects are of the same class while the last is different, e.g., AAAAB.
Attending to Data Contamination.
While data contamination has been explicitly handled in most of the existing benchmarks, object hallucination has been observed even in images that appear in the instruction tuning dataset, such as Visual Genome (zhai2023halle; liu2023mitigating). To evaluate whether multi-object hallucination can be observed in both seen and unseen images, and to critically determine if training on these images helps reduce hallucinations, we explicitly split our dataset into Seen and Unseen based on the original split of the datasets. [^1] We believe this approach is the best practice, but we also acknowledge that the distinction between seen and unseen images may not be strict. Uncurated web images often overlap with public test images, and researchers have no transparent access to the datasets used to train some of these LVLMs unfortunately (dodge2021documenting). Depending on the object query distributions (4 splits) and whether the image appears in the training split (2 splits), we divide the test into 8 folders.
4 Experiments and Results
{threeparttable}
@@ 4 Experiments and Results
[…]
{tablenotes}
Mechanistically grounded LVLMs take visual prompts by dedicated pointer tokens. We slightly adapt the text prompt and probe the object classes with the highest probabilities. We also apply such probabilistic probing to LLaVA-7B for comparison, as all of three models adopt Vicuna-7B v1.5 (Zheng2023JudgingLW) as the base LLM. See Appendix A.1 for details.